OpenStudy (anonymous):

simplify: The sqrt of 150^3

OpenStudy (anonymous):

(2*3*25)^(3/2)=125* 6^(3/2)

OpenStudy (anonymous):

sqrt (150^3) = sqrt (150 * 150 * 150) = sqrt (25* 6 *25* 6*25* 60 = 5 * 5 * 5 sqrt (6 * 6 * 6) = 125 sqrt (36 * 6) = 750 sqrt (6)

OpenStudy (anonymous):

The sqrt of 150b^3. That is my huge mistake, I am so sorry :/

OpenStudy (anonymous):

good job explaining that guys

OpenStudy (anonymous):

\[\sqrt{b^3}=\sqrt{b^2\times b}=\sqrt{b^2}\times \sqrt{b}=b\sqrt{b}\]

OpenStudy (anonymous):

iv had problims with that stuff and i get it a lil beter now

OpenStudy (anonymous):

\[\sqrt{150}=\sqrt{25\times 6}=\sqrt{25}\times\sqrt{6}=5\sqrt{6}\]

OpenStudy (anonymous):

so \[\sqrt{150b^3}=5b\sqrt{6b}\]